# Lorenzo Ciardo

### v1

@@ 3. $(iii)$
As we have discussed in the Overview at the intuitive level, our proof of Theorem 1 works by exploiting the fact that randomly walking over an aperiodic $\mathbf{A}$ for a sufficiently long time makes Alice oblivious of her past locations. Using standard theory of discrete-time Markov chains, [^23] For example, see [38]. we can give this statement a more rigorous meaning. Suppose, for the sake of simplicity, that $\mathbf{A}$ is strongly connected. Alice starts her walk from a vertex $a$ sampled uniformaly at random among all vertices of $\mathbf{A}$. After a number $i$ of time steps, she ascertains her new location $b$ and tries to guess her starting point $a$ based on the known data $b$ and $i$. Clearly, the trivial strategy of choosing at random yields a winning chance of $\frac{1}{n_{\mathbf{A}}}$. Can she do better? Let $T$ be the transition matrix of $\mathbf{A}$; i.e., the row-stochastic matrix $T=D^{-1}M$, where $D$ is the diagonal matrix of the outdegrees of the vertices and $M$ is the adjacency matrix. Since $\mathbf{A}$ is strongly connected, the Markov chain is ergodic, which means that there exists a unique stationary probability distribution $\mathbf{p}$ satisfying $\mathbf{p}^{\top}T=\mathbf{p}^{\top}$. The probability of Alice ending up in $b$ after $i$ time steps starting from $a$ is $\mathbf{e}_{a}^{\top}T^{i}\mathbf{e}_{b}$ (where $\mathbf{e}_{a}$ is the probability distribution assigning all the weight to $a$). By Bayes’ Theorem, to maximise her winning chances, Alice should guess a vertex that maximises the corresponding entry of the vector $T^{i}\mathbf{e}_{b}$ — which is a deterministic vector, as Alice is aware of both $i$ and $b$. Now, it is easy to show that $T$ is primitive if and only if $M$ is primitive. If this is the case, $T^{i}$ converges to the rank-one matrix $\mathbf{1}\mathbf{p}^{\top}$ for $i\to\infty$ (where $\mathbf{1}$ is the all-one vector). Hence, $T^{i}\mathbf{e}_{b}$ converges to the constant vector $p_{b}\mathbf{1}$ — meaning that, after a large-enough number of steps, all choices will roughly yield, for Alice, a disappointing $\frac{1}{n_{\mathbf{A}}}$ winning chance. On the other hand, if the primitivity requirement is relaxed, the Markov chain is not guaranteed to converge to the stationary distribution. In fact, it follows from the Perron–Frobenius Theorem that there exist subsequences of the sequence $(T^{i})_{i\in\mathbb{N}}$ converging to precisely $c$ different limits, where $c$ is the number of eigenvalues of $M$ having maximum modulus [34, § 9.2]; note that $M$ is primitive precisely when $c=1$. In particular, $T^{i}\mathbf{e}_{b}$ may not converge to a constant vector, which makes it possible for Alice to guess correctly with probability strictly larger than $\frac{1}{n_{\mathbf{A}}}$.
@@ 5 Fibrosity
Recall that, given an ($r$-uniform) hypergraph $\mathcal{H}$, $\varphi^{\tau}_{\mathcal{H}}$ denotes the maximum cardinality of a set of mutually disjoint $\tau$-fibers [^24] We point out that, if $f$ and $f^{\prime}$ are two disjoint $\tau$-fibers in $\mathcal{H}$ consisting of the hyperedges $e_{1},\dots,e_{\tau}$ and $e_{1}^{\prime},\dots,e_{\tau}^{\prime}$, respectively, it is not forbidden that, for some $i,j\in[\tau]$, the hyperedges $e_{i}$ and $e_{j}^{\prime}$ are adjacent; however, they cannot be equal. in $\mathcal{H}$, while $\varphi^{\max}_{\mathcal{H}}$, $\pi_{\mathcal{H}}$, and $\lambda_{\mathcal{H}}$ denote the number of maximal fibers, pendent hyperedges, and links in $\mathcal{H}$, respectively. The goal of this section is to prove the following result.
@@
holds, provided that $\mathcal{H}$ has no isolated vertices.
@@
We say that a fiber $f$ connects two (possibly equal) hyperedges $e,e^{\prime}\in E(\mathcal{H})\setminus f$ if some hyperedge in $f$ is adjacent to $e$ and some hyperedge in $f$ is adjacent to $e^{\prime}$. Consider the multigraph (with loops allowed) $\mathcal{G}$ whose vertex set is $\mathscr{Z}\cup\mathscr{W}$ and whose edge multiset is defined as follows: For $e,e^{\prime}\in\mathscr{Z}\cup\mathscr{W}$ (possibly equal), we insert one edge in $\mathcal{G}$ joining $e$ and $e^{\prime}$ for each maximal fiber of $\mathcal{H}$ connecting $e$ and $e^{\prime}$. We claim that $m_{\mathcal{G}}=\varphi^{\max}_{\mathcal{H}}$ (where $m_{\mathcal{G}}$ is the number of edges of the multigraph ${\mathcal{G}}$, counted with multiplicity). First, since any maximal fiber yields at most one edge in ${\mathcal{G}}$, it is clear that $m_{\mathcal{G}}\leq\varphi^{\max}_{\mathcal{H}}$. Suppose that $m_{\mathcal{G}}<\varphi^{\max}_{\mathcal{H}}$. This would mean that there exists a maximal fiber $f$ such that, for each $e\in f$, $e$ is only adjacent to other links. Since $f$ is maximal, $e$ is in fact only adjacent to other hyperedges in $f$. In other words, $f$ induces a cycle, which is impossible as $f$ is non-degenerate by assumption. It follows that the claim is true. For each $e\in\mathscr{Z}\cup\mathscr{W}$, let $\deg_{\mathcal{G}}(e)$ be the degree of $e$ in $\mathcal{G}$; i.e., the number of edges of $\mathcal{G}$ to which $e$ is incident (where a loop contributes $2$). Note that, if $e\in\mathscr{Z}_{i}$ for some $i\in[r]$, $\deg_{\mathcal{G}}(e)=i$. Moreover, if $e\in\mathscr{W}$, $\deg_{\mathcal{G}}(e)=1$. Via the handshaking lemma, we deduce that
@@
We can assume without loss of generality that $\mathcal{H}$ is connected. Let $\mathscr{P}$ and $\mathscr{L}$ denote the set of pendent hyperedges and the set of links of $\mathcal{H}$, respectively. Notice that $\operatorname{sdr}(e)=r-1$ if $e\in\mathscr{L}$, $\operatorname{sdr}(e)\leq r-\frac{7}{6}$ if $e\not\in\mathscr{P}\cup\mathscr{L}$, and, clearly, $\operatorname{sdr}(e)\leq r$ for any hyperedge $e\in E(\mathcal{H})$. Similarly to Proposition 7, the result is obtained by comparing the total $\operatorname{sdr}$ of a certain set of hyperedges to that of all other hyperedges. In this case, instead of the set $\mathscr{Z}$, we consider the set $\mathscr{L}$. Since, by assumption, $\mathcal{H}$ has no isolated vertices, (1) gives
@@ 6 Sparsity
Consider the following parameters:
@@
Notice that all of them are positive real numbers. Let $n$ be an even positive integer number to be determined later, and suppose that $n\geq\ell^{\frac{1}{r-1}}$. Let $p=\ell n^{1-r}$, and sample a hypergraph $\mathcal{H}\sim\mathscr{H}(n,p)$.
@@
Select a set $S$ of vertices of $\mathcal{H}$ of cardinality $\frac{n}{2}$ in a way that each short cycle of $\mathcal{H}$ contains a vertex of $S$, and let $\tilde{\mathcal{H}}$ be the subhypergraph of $\mathcal{H}$ induced by $V(\mathcal{H})\setminus S$. Note that $n_{\tilde{\mathcal{H}}}=\frac{n}{2}$. Furthermore, $\operatorname{ind}(\tilde{\mathcal{H}})\leq\operatorname{ind}(\mathcal{H})<w$, while $\operatorname{gir}(\tilde{\mathcal{H}})\geq g$ since all short cycles have been broken. Since the product of the chromatic and the independence numbers of a hypergraph is at least the number of its vertices (see [9]), we find
@@ 7 Consistency
Similarly, the homonymy $w_{\tau}=x^{(\tau)}_{q_{\tau}}$ does not yield problems as
@@
Furthermore, for $i\in[\tau-1]$, we have $w_{i}=x^{(i+1)}_{p_{i+1}}=x^{(i)}_{q_{i}}$. Also this homonymy is not problematic, as
@@ 8 Everything together
It shall be convenient to only deal with monic structures. The next result shows that this assumption does not yield a loss of generality. While it is essentially folklore, we give a self-contained proof in Appendix A for completeness.
@@
Let $\mathbf{Y}$ be a monic structure whose relation has arity $r$. For $c\in\mathbb{N}$, let $\mathbf{K}_{r,c}$ be the monic structure with domain $[c]$ whose unique, $r$-ary relation is the subset of $[c]^{r}$ obtained by removing all constant tuples. The chromatic number of $\mathbf{Y}$ is the minimum $c$ such that $\mathbf{Y}\to\mathbf{K}_{r,c}$. Note that $\mathbf{Y}$ has a finite chromatic number if and only if it is loopless.
@@
Observe that a structure $\mathbf{Y}$ is aperiodic (resp. loopless) if and only if $\mathbf{Y}^{\operatorname{mon}}$ is aperiodic (resp. loopless). Hence, by virtue of Lemma 28, it is enough to prove the result in the case that $\mathbf{A}$ and $\mathbf{B}$ are monic structures. Let $\tau$ be the mixing time of $\mathbf{A}$ and let $c$ be the chromatic number of $\mathbf{B}$; since $\mathbf{A}$ is aperiodic and $\mathbf{B}$ is loopless, these numbers are both finite. Also, let $R$ be the unique symbol in their signature and let $r$ be the arity of $R$. If $r=1$ or if $r=2$ and $\tau=1$, Definition 3 (and Footnote 15) would imply that $\mathbf{A}$ contains a loop, which is impossible since $\mathbf{A}\to\mathbf{B}$ and $\mathbf{B}$ is loopless.
@@
Take $n\in\mathbb{N}$. We can assume without loss of generality that $n\geq\frac{1}{\epsilon}$. Otherwise, we might simply take as $\mathbf{X}$ any instance on $n$ vertices having a loop. Since $\mathbf{B}$ is loopless, $\mathbf{X}\not\to\mathbf{B}$. Moreover, $\operatorname{LC}^{\lfloor\epsilon n\rfloor}(\mathbf{X},\mathbf{A})=\operatorname{LC}^{0}(\mathbf{X},\mathbf{A})=\textsc{Yes}$, as needed. The fact that $n\geq\frac{1}{\epsilon}$ implies that $n\geq n_{0}$, via (13). Hence, invoking Theorem 10, we find a $(\delta n,\beta)$-threshold-sparse $r$-uniform hypergraph $\mathcal{H}$ such that $n_{\mathcal{H}}=n$, $\operatorname{gir}(\mathcal{H})\geq\tau$, and $\operatorname{chr}(\mathcal{H})\geq c+1$. Consider now a structure $\mathbf{X}$ obtained from $\mathcal{H}$ by choosing an arbitrary orientation for each hyperedge; i.e., $\mathbf{X}$ has domain $V(\mathcal{H})$ and its unique relation $R^{\mathbf{X}}$ (of arity $r$) contains, for each $e\in E(\mathcal{H})$, precisely one $r$-tuple $\mathbf{x}$ such that $\operatorname{set}(\mathbf{x})=e$. Note that $\mathbf{X}$ is an oriented monic structure, and ${\mathbf{X}^{\operatorname{sym}}}=\mathcal{H}$. We claim that the pair $(\mathbf{X},\mathbf{A})$ has a $(\lfloor\epsilon n\rfloor,\delta n)$-consistency gap. Take a homomorphism $f:\mathbf{W}\to\mathbf{A}$, where $\mathbf{W}$ is a substructure of $\mathbf{X}$ with $n_{\mathbf{W}}\leq\lfloor\epsilon n\rfloor$. Suppose, for the sake of contradiction, that there exists a substructure $\mathbf{Y}$ of $\mathbf{X}$ such that $f$ is not consistent with $\mathbf{Y}$ and $\frac{\delta n}{n_{\mathbf{A}}}<m_{\mathbf{Y}}\leq\delta n$, and assume without loss of generality that $\mathbf{Y}$ is minimal, in the sense that $f$ is consistent with any proper substructure of $\mathbf{Y}$. Let $\mathcal{G}={\mathbf{Y}^{\operatorname{sym}}}$, and let $\mathcal{G}^{\prime}$ be the subhypergraph of $\mathcal{G}$ obtained by removing all isolated vertices of $\mathcal{G}$. By the monotonicity of threshold-sparsity, we deduce that $\mathcal{G}^{\prime}$ is hereditarily $\beta$-sparse. Moreover, its girth is at least $\tau$. It follows from Theorem 6 that

### v2

@@ 3. $(iii)$
As we have discussed in the Overview at the intuitive level, our proof of Theorem 1 works by exploiting the fact that randomly walking over an aperiodic $\mathbf{A}$ for a sufficiently long time makes Alice oblivious of her past locations. Using standard theory of discrete-time Markov chains, [^16] For example, see [42]. we can give this statement a more rigorous meaning. Suppose, for the sake of simplicity, that $\mathbf{A}$ is strongly connected. Alice starts her walk from a vertex $a$ sampled uniformaly at random among all vertices of $\mathbf{A}$. After a number $i$ of time steps, she ascertains her new location $b$ and tries to guess her starting point $a$ based on the known data $b$ and $i$. Clearly, the trivial strategy of choosing at random yields a winning chance of $\frac{1}{n_{\mathbf{A}}}$. Can she do better? Let $T$ be the transition matrix of $\mathbf{A}$; i.e., the row-stochastic matrix $T=D^{-1}M$, where $D$ is the diagonal matrix of the outdegrees of the vertices and $M$ is the adjacency matrix. Since $\mathbf{A}$ is strongly connected, the Markov chain is ergodic, which means that there exists a unique stationary probability distribution $\mathbf{p}$ satisfying $\mathbf{p}^{\top}T=\mathbf{p}^{\top}$. The probability of Alice ending up in $b$ after $i$ time steps starting from $a$ is $\mathbf{e}_{a}^{\top}T^{i}\mathbf{e}_{b}$ (where $\mathbf{e}_{a}$ is the probability distribution assigning all the weight to $a$). By Bayes’ Theorem, to maximise her winning chances, Alice should guess a vertex that maximises the corresponding entry of the vector $T^{i}\mathbf{e}_{b}$—which is a deterministic vector, as Alice is aware of both $i$ and $b$. Now, it is easy to show that $T$ is primitive if and only if $M$ is primitive. If this is the case, $T^{i}$ converges to the rank-one matrix $\mathbf{1}\mathbf{p}^{\top}$ for $i\to\infty$ (where $\mathbf{1}$ is the all-one vector). Hence, $T^{i}\mathbf{e}_{b}$ converges to the constant vector $p_{b}\mathbf{1}$—meaning that, after a large-enough number of steps, all choices will roughly yield, for Alice, a disappointing $\frac{1}{n_{\mathbf{A}}}$ winning chance. On the other hand, if the primitivity requirement is relaxed, the Markov chain is not guaranteed to converge to the stationary distribution. In fact, it follows from the Perron–Frobenius Theorem that there exist subsequences of the sequence $(T^{i})_{i\in\mathbb{N}}$ converging to precisely $c$ different limits, where $c$ is the number of eigenvalues of $M$ having maximum modulus [38, § 9.2]; note that $M$ is primitive precisely when $c=1$. In particular, $T^{i}\mathbf{e}_{b}$ may not converge to a constant vector, which makes it possible for Alice to guess correctly with probability strictly larger than $\frac{1}{n_{\mathbf{A}}}$.
@@ 5 Fibrosity
Recall that, given an ($r$-uniform) hypergraph $\mathcal{H}$, $\varphi^{\tau}_{\mathcal{H}}$ denotes the maximum cardinality of a set of mutually disjoint $\tau$-fibers in $\mathcal{H}$, while $\varphi^{\max}_{\mathcal{H}}$, $\pi_{\mathcal{H}}$, and $\lambda_{\mathcal{H}}$ denote the number of maximal fibers, pendent hyperedges, and links in $\mathcal{H}$, respectively. The goal of this section is to prove the following result.
@@
holds, provided that $\mathcal{H}$ has no isolated vertices. We point out that, if $f$ and $f^{\prime}$ are two disjoint $\tau$-fibers in $\mathcal{H}$ consisting of the hyperedges $e_{1},\dots,e_{\tau}$ and $e_{1}^{\prime},\dots,e_{\tau}^{\prime}$, respectively, it is not forbidden that, for some $i,j\in[\tau]$, the hyperedges $e_{i}$ and $e_{j}^{\prime}$ are adjacent; however, they cannot be equal.
@@
We say that a fiber $f$ connects two (possibly equal) hyperedges $e,e^{\prime}\in E(\mathcal{H})\setminus f$ if some hyperedge in $f$ is adjacent to $e$ and some hyperedge in $f$ is adjacent to $e^{\prime}$. Consider the multigraph (with loops allowed) $\mathcal{G}$ whose vertex set is $\mathscr{Z}\cup\mathscr{W}$ and whose edge multiset is defined as follows: For $e,e^{\prime}\in\mathscr{Z}\cup\mathscr{W}$ (possibly equal), we insert one edge in $\mathcal{G}$ joining $e$ and $e^{\prime}$ for each maximal fiber of $\mathcal{H}$ connecting $e$ and $e^{\prime}$. We claim that $m_{\mathcal{G}}=\varphi^{\max}_{\mathcal{H}}$ (where $m_{\mathcal{G}}$ is the number of edges of the multigraph ${\mathcal{G}}$, counted with multiplicity). First, since any maximal fiber yields at most one edge in ${\mathcal{G}}$, it is clear that $m_{\mathcal{G}}\leq\varphi^{\max}_{\mathcal{H}}$. Suppose that $m_{\mathcal{G}}<\varphi^{\max}_{\mathcal{H}}$. This would mean that there exists a maximal fiber $f$ such that, for each $e\in f$, $e$ is only adjacent to other links. Since $f$ is maximal, $e$ is in fact only adjacent to other hyperedges in $f$. In other words, $f$ induces a cycle, which is impossible as $f$ is non-degenerate by assumption. It follows that the claim is true. For each $e\in\mathscr{Z}\cup\mathscr{W}$, let $\deg_{\mathcal{G}}(e)$ be the degree of $e$ in $\mathcal{G}$; i.e., the number of edges of $\mathcal{G}$ to which $e$ is incident (where a loop contributes $2$). Note that, if $e\in\mathscr{Z}_{i}$ for some $i\in[r]$, $\deg_{\mathcal{G}}(e)=i$. Moreover, if $e\in\mathscr{W}$, $\deg_{\mathcal{G}}(e)=1$. We can now express the number of edges in $\mathcal{G}$ in terms of the sum of the vertex degrees via the handshaking lemma, thereby deducing that
@@
We point out that, in the proof of Proposition 7, we are crucially using that the upper bound $\operatorname{sdr}(e)\leq r-\frac{i}{2}$ on the sum of the degree reciprocals of a hyperedge $e\in\mathscr{Z}_{i}$ can be improved in the case that $i=1$ or $i=2$. Without the improvement, the resulting upper bound on the number of maximal fibers of $\mathcal{H}$ would be $\varphi^{\max}_{\mathcal{H}}<(\beta-1)n_{\mathcal{H}}+\pi_{\mathcal{H}}+z$, and the presence of $z$ in the new expression would make the proposition useless towards proving Theorem 6.
@@
Like in the proof of Proposition 7, we can assume without loss of generality that $\mathcal{H}$ is connected. Let $\mathscr{P}$ and $\mathscr{L}$ denote the set of pendent hyperedges and the set of links of $\mathcal{H}$, respectively. Similarly to Proposition 7, the result is obtained by comparing the total $\operatorname{sdr}$ of a certain set of hyperedges to that of all other hyperedges. In this case, instead of the set $\mathscr{Z}$, we consider the set $\mathscr{L}$.
Notice that $\operatorname{sdr}(e)\leq r$ for any hyperedge $e\in E(\mathcal{H})$ and $\operatorname{sdr}(e)=r-1$ if $e\in\mathscr{L}$. Furthermore, if $e\not\in\mathscr{P}\cup\mathscr{L}$, there are two possibilities: Either $(i)$ three or more vertices of $e$ have degree at least two, or $(ii)$ one vertex of $e$ has degree at least three, one other vertex has degree at least two, and all remaining vertices have degree one; in both cases, we find $\operatorname{sdr}(e)\leq r-\frac{7}{6}$. Since, by assumption, $\mathcal{H}$ has no isolated vertices, (1) gives
@@ 6 Sparsity
We start by defining five parameters $\ell$, $\nu$, $\vartheta$, $\mu$, and $\delta$. Intuitively, $\ell$ determines the hyperedge probability of the Erdős-Rényi hypergraph we are soon going to sample; $\nu$ and $\mu$ express the desired vertex-threshold-sparsity of the hypergraph; $\delta$ controls the threshold-sparsity of a subhypergraph we shall obtain by removing all short cycles; and $\vartheta$ simply connects the other parameters in a way that all necessary inequalities in the proof hold.
@@
Notice that all of the parameters are positive real numbers. Let $n$ be an even positive integer number to be determined later, and suppose that $n\geq\ell^{\frac{1}{r-1}}$. Let $p=\ell n^{1-r}$, and sample a hypergraph $\mathcal{H}\sim\mathscr{H}(n,p)$.
@@
Select a set $S$ of vertices of $\mathcal{H}$ of cardinality $\frac{n}{2}$ in a way that each short cycle of $\mathcal{H}$ contains a vertex of $S$; in order for its cardinality to be exactly $\frac{n}{2}$, $S$ is allowed to contain vertices that do not appear in any short cycle. Let $\tilde{\mathcal{H}}$ be the subhypergraph of $\mathcal{H}$ induced by $V(\mathcal{H})\setminus S$, so that $n_{\tilde{\mathcal{H}}}=\frac{n}{2}$. We have $\operatorname{ind}(\tilde{\mathcal{H}})\leq\operatorname{ind}(\mathcal{H})<w$, while $\operatorname{gir}(\tilde{\mathcal{H}})\geq g$ as all short cycles have been broken. Since the product of the chromatic and the independence numbers of a hypergraph is at least the number of its vertices (see [9]), we find
@@ 7 Consistency
Similarly, the fact that $w_{\tau}=x^{(\tau)}_{q_{\tau}}$ does not yield problems as
@@
Furthermore, for $i\in[\tau-1]$, we have $w_{i}=x^{(i+1)}_{p_{i+1}}=x^{(i)}_{q_{i}}$. Also this identity is not problematic, as
@@ 8 Everything together
It shall be convenient to only deal with monic structures. The next result shows that this assumption does not yield a loss of generality. While it is essentially folklore, we give a self-contained proof in the Appendix for completeness.
@@
Let $\mathbf{Y}$ be a monic structure whose relation has arity $r$. For $c\in\mathbb{N}$, let $\mathbf{K}_{r,c}$ be the monic structure with domain $[c]$ whose unique, $r$-ary relation is the subset of $[c]^{r}$ obtained by removing all constant tuples. If $\mathbf{Y}$ is loopless, its chromatic number is the minimum $c$ such that $\mathbf{Y}\to\mathbf{K}_{r,c}$.
@@
Observe that a structure $\mathbf{Y}$ is aperiodic (resp. loopless) if and only if $\mathbf{Y}^{\operatorname{mon}}$ is aperiodic (resp. loopless). Hence, by virtue of Lemma 28, it is enough to prove the result in the case that $\mathbf{A}$ and $\mathbf{B}$ are monic structures. Let $\tau$ be the mixing time of $\mathbf{A}$ and let $c$ be the chromatic number of $\mathbf{B}$; since $\mathbf{A}$ is aperiodic and $\mathbf{B}$ is loopless, these numbers are both finite. Also, let $R$ be the unique symbol in their signature and let $r$ be the arity of $R$. If $r=1$ or if $r=2$ and $\tau=1$, Definition 3 (and Footnote 11) would imply that $\mathbf{A}$ contains a loop, which is impossible since $\mathbf{A}\to\mathbf{B}$ and $\mathbf{B}$ is loopless. Thus, we can assume that either $r\geq 3$ or $\tau\geq 2$ (as in the proof of Theorem 14).
@@
Take $n\in\mathbb{N}$. We can assume without loss of generality that $n\geq\frac{1}{\epsilon}$. Otherwise, we might simply take as $\mathbf{X}$ any instance on $n$ vertices having a loop. Since $\mathbf{B}$ is loopless, $\mathbf{X}\not\to\mathbf{B}$. Moreover, $\operatorname{LC}^{\lfloor\epsilon n\rfloor}(\mathbf{X},\mathbf{A})=\operatorname{LC}^{0}(\mathbf{X},\mathbf{A})=\textsc{Yes}$, as needed. The fact that $n\geq\frac{1}{\epsilon}$ implies that $n\geq n_{0}$, via (13). Hence, invoking Theorem 10, we find a $(\delta n,\beta)$-threshold-sparse $r$-uniform hypergraph $\mathcal{H}$ such that $n_{\mathcal{H}}=n$, $\operatorname{gir}(\mathcal{H})\geq\tau$, and $\operatorname{chr}(\mathcal{H})\geq c+1$. Consider now a structure $\mathbf{X}$ obtained from $\mathcal{H}$ by choosing an arbitrary orientation for each hyperedge; i.e., $\mathbf{X}$ has domain $V(\mathcal{H})$ and its unique relation $R^{\mathbf{X}}$ (of arity $r$) contains, for each $e\in E(\mathcal{H})$, precisely one $r$-tuple $\mathbf{x}$ such that $\operatorname{set}(\mathbf{x})=e$. Note that $\mathbf{X}$ is an oriented monic structure, and ${\mathbf{X}^{\operatorname{sym}}}=\mathcal{H}$. We claim that the pair $(\mathbf{X},\mathbf{A})$ has a $(\lfloor\epsilon n\rfloor,\delta n)$-consistency gap. Take a homomorphism $f:\mathbf{W}\to\mathbf{A}$, where $\mathbf{W}$ is a substructure of $\mathbf{X}$ with $n_{\mathbf{W}}\leq\lfloor\epsilon n\rfloor$. Suppose, for the sake of contradiction, that there exists a substructure $\mathbf{Y}$ of $\mathbf{X}$ such that $f$ is not consistent with $\mathbf{Y}$ and $\frac{\delta n}{n_{\mathbf{A}}}<m_{\mathbf{Y}}\leq\delta n$. Assume, without loss of generality, that $\mathbf{Y}$ is minimal, in the sense that $f$ is consistent with any proper substructure of $\mathbf{Y}$. Let $\mathcal{G}={\mathbf{Y}^{\operatorname{sym}}}$, and let $\mathcal{G}^{\prime}$ be the subhypergraph of $\mathcal{G}$ obtained by removing all isolated vertices of $\mathcal{G}$. By the monotonicity of threshold-sparsity, we deduce that $\mathcal{G}^{\prime}$ is hereditarily $\beta$-sparse. Moreover, its girth is at least $\tau$. It follows from Theorem 6 that